Current Trends in Communications Industry and Regulations

Gregg Sayre

Rochester Telephone

Abstract

The telecommunications industry is in a turmoil, faced with massive structural changes likely to result from pending Federal legislation; accelerating technological improvements, sparking competition across all sectors of the industry; and internal restructuring, mergers and downsizing to try to make a success in the new environment. No one can credibly predict the future of this industry over more than a few years, but I think a few trend lines can be drawn. I will discuss in detail the current emergence of competition in the Frontier/Rochester Telephone Open Market Plan and make some predictions about the impact of regulatory change, technological progress and competition on the key players in the industry, with a focus on local telephone companies and their customers. My thesis is that the big fish will eat the little fish, but enough fish will survive to provide more and better services at lower prices.

Coffee and cookies will be served